OpenStudy (anonymous):

HELP! During the period 1998–2002, the number y (in millions) of juvenile books shipped to bookstores can be modeled by the equation y = -15x^2 + 64x + 360 where x is the number of years since 1998. During what year were there 400 million juvenile books shipped to bookstores?

OpenStudy (anonymous):

y = -15x^2 + 64x + 360 400 = -15x^2 + 64x + 360 -15x^2 + 64x - 40 = 0 sol ve this for x

OpenStudy (anonymous):

- a bit messy this one!

OpenStudy (anonymous):

i think this is best solved using the quadratic formula

OpenStudy (anonymous):

Year of 2001?

OpenStudy (anonymous):

400=-15x^2 + 64x + 360

OpenStudy (anonymous):

if year is 2001, ... y=-15(2001)^2 + 64(2001) + 360

OpenStudy (anonymous):

x = year, ok?

OpenStudy (anonymous):

hold on - solving x = 4.82 so that must be 1998 + 5 = 2003

OpenStudy (anonymous):

ooopps - i made an error solvig

OpenStudy (anonymous):

i got my signs wrong correct is x = 3.51 or 0.76

OpenStudy (anonymous):

Which would be 1998 + 3.51? = 2001.51 or 2001?

OpenStudy (anonymous):

- yes thats a good question - i think it must be 2001 because 2001 . 51 is half way through that year

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

yw - good luck